OpenStudy (aperez):

Find the distance between points P(2, 2) and Q(7, 4) to the nearest tenth.

OpenStudy (anonymous):

5.4 rounded

OpenStudy (anonymous):

D={(7-2)^2+(4-2)^2}^0.5 = 29^0.5 =5.4

OpenStudy (anonymous):

\[d = \sqrt{(x2-x1)^{2}+(y2-y1)^{2}}\]

OpenStudy (aperez):

Thankyou very much.

OpenStudy (anonymous):

Close the question